27 May 2009

We got a roof on, and the weather begins to act funny

A scene from laying roof on the elevator trench.

A scene from laying roof on the elevator trench.

Sarah prepared delicious musk-ox for dinner.

What we have done today:

  1. Drilling. We got a record core length of 2.77 m
  2. Logging ice cores. Last bag logged: 308, depth 169.40 m.
  3. All markers for the new skiway are up. Everything is groomed with tiller. Only a faint trace in the snow of the old skiway remains.
  4. Excavating for elevator and staircase.
  5. Building up in the CFA laboratory.
  6. Elevator installed in trench with crane.
  7. Setting up tables and saws in science trench. Now only a few tables remain to be set up.
  8. Putting roof on elevator shaft and stairwell and trench. The crew worked until 23.00.
  9. Rearranging 2nd and 3rd floor in main dome.

Ad.1: This is how the drillers' report looks today:
“Close inspection found that one of the three screws that hold the inner pump sleeve fast to the outer barrel was missing and presumed to be in the borehole. It was a good candidate for the troubled runs that closed out yesterday’s drilling. The magnet was mounted to the drill head to fish it out. The magnet run surfaced without a screw, but at the same time Nobby found the small screw in the chips from the final run brought up the previous night. With a clean borehole the drilling continued with the EPICA drill. The lower valve was remounted to the pump and the hollow shaft opened resulting in a return to the good mode of drilling we experienced before this valve was removed. A modification made to the pump at NGRIP that closed the hollow shaft to the surrounding environment within the superbanger had not been installed on the pump installed on the EPICA drill. The pump/hollow shaft configuration we have now works well with this drill liquid because any excess chips in the borehole do not accumulate at the bottom but stay suspended or float to the top of the liquid column.
Driller’s depth is 158.10 m.”

Ad.9: We have assembled sofas and have made a lounge with sofa group and tables on the 2nd floor. The Field Leader office has been moved to the 3rd floor.

Weather: Clear in the morning, at noon overcast and snow, some clearing in the evening but still snow, -23°C to -11°C, calm-10 knots from NE, later turning via S to NW.
Visibility: 500 m to 5 km. Snow.

FL, J.P. Steffensen

← Previous entry   Next entry →